if I'm told the area of a triangle is 81 inches squared and the base is twice the height how do I solve?
Answer by rfer(16322)

2x*x=81
2x^2=81
x^2=81/2
x^2=40.5
sqrt 40.5=6.36 in
2x=12.73 in